# Minds versus Machines: Rethinking Entailment Verification with Language Models

### v1

@@ Abstract
Humans make numerous inferences in text comprehension to understand discourse. This paper aims to understand the commonalities and disparities in the inference judgments between humans and state-of-the-art Large Language Models (LLMs). Leveraging a comprehensively curated entailment verification benchmark, we evaluate both human and LLM performance across various reasoning categories. Our benchmark includes datasets from three categories (NLI, contextual QA, and rationales) that include multi-sentence premises and different knowledge types, thereby evaluating the inference capabilities in complex reasoning instances. Notably, our findings reveal LLMs’ superiority in multi-hop reasoning across extended contexts, while humans excel in tasks necessitating simple deductive reasoning. Leveraging these insights, we introduce a fine-tuned Flan-T5 model [^1] https://huggingface.co/soumyasanyal/entailment-verifier-xxl that outperforms GPT-3.5 and rivals with GPT-4, offering a robust open-source solution for entailment verification. As a practical application, we showcase the efficacy of our finetuned model in enhancing self-consistency in model-generated explanations, resulting in a 6% performance boost on average across three multiple-choice question-answering datasets.
Minds versus Machines: Rethinking Entailment Verification with Language Models
@@ 1 Introduction
A prevailing notion in cognitive psychology exists that humans make numerous inferences to understand discourse and text Garnham (1989). These inferences, serving diverse functions in text comprehension, extend beyond resolving lexical ambiguities or determining the referent of a pronoun; they play a crucial role in linking information from disparate sections of a text to establish its literal meaning. Cognitive studies Buschman et al. (2011); Cowan (2001) have shown that an average human brain has a limited capacity to retain only four chunks in short-term memory, indicating a limitation of human inference abilities. Similarly, prior works Gururangan et al. (2018); McCoy et al. (2019); Geiger et al. (2020); Clark et al. (2020); Sanyal et al. (2022) have shown that large language models (LLMs) are still subpar at understanding negations in inference, use spurious correlations while reasoning, cannot reason through multi-step compositions robustly, etc. With the advent of large language models (LLMs) and their applications on natural language inference (NLI) Minaee et al. (2021); Nie et al. (2020), understanding the key difference in the inference abilities between humans and LLMs is crucial to making further progress, which is currently missing.
In this work, we aim to address this by first evaluating both humans and current LLMs Liu et al. (2019); Tafjord et al. (2022); Chung et al. (2022); Brown et al. (2020); OpenAI (2023) on a comprehensive entailment verification benchmark and then analyzing the performance across different reasoning categories based on the ease of inference deduction and the type of knowledge required. Existing textual inference datasets such as SNLI Bowman et al. (2015), MNLI Williams et al. (2018), etc., mostly contain short sentence premises that do not completely encapsulate the challenges of multi-sentence, complex reasoning. Predicting the entailment of such complex premise-hypothesis pairs often requires multi-hop reasoning, inferring missing information, robustness to spurious correlation, etc., which are largely missing from simpler datasets Gururangan et al. (2018); McCoy et al. (2019). Also, real-world scenarios that require inference skills, such as understanding stories, dialogues, etc., usually contain multi-sentence premises. Therefore, we select multiple datasets across three categories (NLI, contextual QA, and rationales) and convert them into NLI format, as required, to create the evaluation benchmark. As shown in Table LABEL:tab:dataset_comparisons, the datasets used in our study typically contain multi-sentence premises that require different types of knowledge to predict the entailment.
@@
Our analyses show that LLMs are stronger than humans at entailment tasks that involve multi-hop reasoning across long contexts. This suggests that training LLMs to reason on diverse long context data is beneficial for entailment verification on complex cases. In contrast, we find that humans are better at solving entailment tasks that require simple deductive reasoning using substitutions, negations, etc., indicating that current LLMs still lack consistency along these reasoning aspects. Further, we find that humans and LLMs perform comparably in entailment tasks requiring inferring some missing knowledge. These findings are further depicted in Figure LABEL:fig:motivation with motivating examples.
@@
Among the different LLMs, we find that general instruction-finetuned models are better than task-finetuned models trained on a specific dataset category. We leverage this finding by finetuning a Flan-T5 Chung et al. (2022) model on a training subset containing datasets from each category. We explore two different training approaches: a classification-based finetuning that learns to directly predict the label and a ranking-based finetuning that learns to rank the most supported hypothesis from a given pair of hypotheses for a given premise. We find that ranking-based finetuning is superior to classification as it can learn a softer decision boundary. Overall, our fine-tuned models outperform GPT-3.5 and perform comparably to GPT-4 on the benchmark, thus providing a strong open-sourced model for entailment verification.
@@
Finally, we demonstrate the utility of our finetuned models on a downstream application of filtering unfaithful model-generated explanations. Recent works Wei et al. (2022); Zhou et al. (2023); Wang et al. (2023) have prompted LLMs to generate natural language explanations for model predictions, but such reasoning steps can sometimes be inconsistent, i.e., the explanation does not entail the model’s final prediction Ye and Durrett (2022). In self-consistency (SC) Wang et al. (2023), multiple reasoning paths are first sampled from the LLM decoder for a given instance and then aggregated to predict the most consistent answer. We use our finetuned models to filter out non-entailed reasoning chains in the SC decoding strategy before aggregating the final prediction, which leads to 6% performance improvement on average across three MCQ datasets.
2 Humans vs. LLMs on Entailment Verification
@@
In this section, we define the task of entailment verification (EV), datasets used to create the evaluation benchmark, the evaluation procedure for evaluating different LLM baselines and humans, and compare them across different reasoning types.
[…]
@@ 2.1 Evaluation Benchmark
To benchmark humans and LLMs on the task of entailment verification, a basic criterion for dataset selection is that the dataset should contain a premise, a hypothesis, and a support/not support label associated with it. Additionally, we list some more desirable characteristics we want to cover in this study, as discussed below:
@@
  • Type of Knowledge: Often, one or more information in the premise needs to be used to predict support. We categorize this information as entity-grounded, commonsense, or localized. Entity-grounded knowledge consists of information about entities and other general knowledge verifiable on the internet. These can be facts about general science, history, etc., or details of some known person, event, etc. It is possible to infer this information even if not mentioned in the premise. The commonsense knowledge is typically all information about everyday life that humans use implicitly but cannot always be verified online. This information is often missing from the premise and has to be inferred implicitly. Lastly, localized information is all other knowledge provided for understanding the events, people, or items mentioned in the premise that is not grounded to any known entity. This information depends on the premise’s specific context and, thus, is impossible to infer unless stated explicitly. Please refer to Table LABEL:tab:knowledge_examples for examples of each knowledge type.
@@
Natural Language Inference
NLI is an obvious choice of data source. While converting an NLI dataset for our task, we merge the neutral and contradict labels to the not support label. We use the following NLI datasets in our benchmark: WaNLI Liu et al. (2022), FEVER Nie et al. (2019), and ANLI Nie et al. (2020).
@@
Next, we consider QA datasets where the task is to answer a question based on a given context and some options. We use an off-the-shelf QA-to-statement converter model Chen et al. (2021) to generate a hypothesis statement for each question option pair. Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support” to create the entailment verification dataset. Overall, we include the following datasets from this category: Cosmos QA (CosQA) Huang et al. (2019), SocialIQA (SIQA) Sap et al. (2019), DREAM Sun et al. (2019), BoolQ Clark et al. (2019), and RACE Lai et al. (2017).
@@
Lastly, we consider data sources where human-annotated explanations are available that justify the original hypothesis (or the correct option, in the case of QA datasets). In this case, we use the rationales as the premise. We use the following datasets: Entailer Tafjord et al. (2022), and ECQA Aggarwal et al. (2021).
@@ 2.2 LLM Evaluation Setup
In this category, the models considered are already finetuned for either NLI or the exact entailment verification task itself. We evaluate RoBERTa (finetuned on NLI data) and Entailer-11B (finetuned on Entailer data) in this category. Please refer to Appendix B.1 for more details on the evaluation setup for these models.
Instruction-finetuned LLMs
These are “general-purpose” language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions. Here, we evaluate Flan-T5-xxl, GPT-3.5, and GPT-4 models. To compute the label, we first modify a given premise-hypothesis pair $(p,h)$ into a prompted input $\mathcal{P}$ using the prompt template as shown in Box 1. Next, we compute a score $s$ as defined below:
where $p_{LLM}(\cdot|\mathcal{P})$ is the model’s probability distribution over the vocabulary. If the score $s$ is higher than a threshold (typically set to $0.5$ in all our experiments), we assign the label support, else we assign the label not support. For GPT-4 evaluation, we directly check for the “Yes” / “No” label prediction as the token probabilities are not accessible via the API. Please refer to Appendix B.2 for more details about the models and ablations on few-shot prompts.
@@
2.3 Evaluation Metric
2.4 Human vs. LLMs
First, we randomly sample 100 instances for each dataset (i.e., 1000 instances in total) and conduct a human evaluation on this subset to estimate average human performance. Please refer to Appendix E.1 for more details on the annotation procedure. Additionally, we evaluate the above LLMs on this sampled subset and report those numbers for fair comparisons with humans. Table LABEL:tab:human_results shows the overall evaluation results. Among LLMs, we observe that instruction-finetuned LLMs are stronger than task-finetuned models that are trained on some selected datasets. This shows that using a specific dataset for fine-tuning does not necessarily improve the general inference abilities of LLMs. On comparisons between humans and LLMs, we find that humans beat all the baseline LLMs, except GPT-4. This shows that existing open-sourced LLMs are not at par with humans on this task. Additionally, although humans and GPT-4 perform comparably on average, there are large misalignments in different individual datasets. Specifically, we observe that ANLI, CosQA, SIQA, and Entailer are the four datasets with $>0.1$ absolute macro-F1 difference. Next, we perform more analysis on these four datasets to better understand these misalignments.
@@ 2.3 Evaluation Metric
2.5 Reasoning Type Analysis
We design a new analysis to further understand the reasons behind such misalignment between models and humans. First, we categorize the type of reasoning required to successfully predict an entailment into the following four categories:
  • Simple Deductive (R1): In this type, the premise contains sentences that can be minimally combined in one step to predict the support for the hypothesis. This type typically tests skills such as substitution, negation, synonyms, etc.
  • Complex Deductive (R2): In this, more than one step of reasoning using information from the premise is required to solve the task. Typically, this type tests skills like mathematical reasoning, using multiple information in context to arrive at a conclusion, etc.
  • Missing Entity-grounded/Commonsense Knowledge (R3): This is a case where some essential commonsense or entity-grounded knowledge is missing in the premise. This gap can potentially be bridged by the model’s parametric knowledge obtained from pretraining.
  • Missing Localized Knowledge (R4): In this, information very specific to the context of the premise is missing. Typically, this is information about the subjects in the context and is not grounded to any entities that can be known via the internet. It is practically impossible for the model (or even humans) to infer such missing information.
Please refer to Appendix E.3 for examples of each reasoning type. Given this categorization, we first annotate the reasoning type of the 100 previously sampled instances for each dataset with absolute macro-F1 difference $>0.1$. We make this choice intending to attribute the largely misaligned datasets (namely, ANLI, CosQA, SIQA, and Entailer) since some random noise in the annotation and sampling process can potentially also cause some misalignment. Please refer to Appendix E.2 for more details on the annotation setup. Figure LABEL:fig:reasoning_analysis2 depicts the aggregated results for each reasoning type. The top plot shows the frequency of each reasoning type, and the bottom plot compares the human and GPT-4 macro-F1 scores. Please refer to Appendix E.4 for a detailed analysis of individual datasets.
The first type in Figure LABEL:fig:reasoning_analysis2 is simple deductive reasoning ($\sim$ 25% data). Here, humans perform better than GPT-4 by a small margin. Instances that require simple deductive reasoning usually use substitutions, negations, paraphrasing, etc., to prove entailment (refer to Table LABEL:tab:reasoning_examples for examples). We find that humans are more robust than GPT-4 in performing such simple deductive reasoning tasks, as observed in prior works Sanyal et al. (2022); Nguyen et al. (2023).
@@ 2.5 Reasoning Type Analysis
Next, we find that GPT-4 significantly outperforms the human baseline on complex reasoning that constitutes $\sim$ 20% data. This type of reasoning usually requires two skills: understanding multiple relevant information in the premise and combining them for reasoning. There are two potential reasons for this misalignment. First, compared to humans, GPT-4 is likely a stronger context processor, especially for long premises, since it has been trained on long-context data sources OpenAI (2023). Alternatively, another potential confounding factor is that AMT workers are typically incentivized to quickly finish annotations to maximize earnings. Thus, instances that require multi-step reasoning can have noisier annotations, potentially underestimating true human performance.
Lastly, we observe that approximately 30% of the data has some missing entity-grounded or commonsense information while $\sim$ 25% of the data has missing localized information. To correctly predict entailment in such instances, a system should be able to infer some of the missing grounded knowledge while not hallucinating specific localized information not mentioned in the premise. We find that both humans and models are comparable across the reasoning types R3 and R4, with R4 being more challenging. This shows that both models and humans tend to hallucinate missing localized information.

### v3

# Are Machines Better at Complex Reasoning? Unveiling Human-Machine Inference Gaps in Entailment Verification

@@ Abstract
Making inferences in text comprehension to understand the meaning is essential in language processing. This work studies the entailment verification (EV) problem of multi-sentence premises that requires a system to make multiple inferences implicitly. Studying EV for such complex premises is important because modern NLP problems, such as detecting inconsistent model-generated rationales, require complex multi-hop reasoning. However, current textual inference datasets mostly contain short premises that only partially focus on these challenges. To address this, we compile an EV benchmark that includes datasets from three NLP domains (NLI, contextual QA, and rationales) containing multi-sentence premises. On benchmarking humans and LLMs, we find that LLMs are better than humans in multi-hop reasoning across extended contexts, while humans perform better in simple deductive reasoning tasks. We also finetune a Flan-T5 model [^1] https://huggingface.co/soumyasanyal/entailment-verifier-xxl for EV using two training objectives to obtain a strong open-source model that outperforms GPT-3.5 and rivals GPT-4. Finally, we use this model to filter out inconsistent model-generated rationales in self-consistency decoding, resulting in a 6% accuracy improvement on average across three MCQ datasets.
Are Machines Better at Complex Reasoning? Unveiling Human-Machine Inference Gaps in Entailment Verification
@@ 1 Introduction
A prevailing notion in cognitive psychology exists that humans make numerous inferences to understand discourse and text Garnham (1989). These inferences play a crucial role in linking information from disparate sections of a text to establish its literal meaning and are closely associated with reasoning. With the recent applications of Large Language Models (LLMs) Devlin et al. (2019); Raffel et al. (2020); Brown et al. (2020); Chung et al. (2022) in NLP tasks that require inference skills Minaee et al. (2021), it is thus essential to understand and improve upon the limitations of LLMs concerning different aspects of language inferences.
In this work, we focus on the task of entailment verification (EV) that classifies whether a given context supports a hypothesis. To assert the validity of a hypothesis, a system has to make multiple inferences from the given context and its internal knowledge, which requires complex multi-hop reasoning. [^2] In a multi-hop reasoning instance, a system has to infer implicit inferences by combining information from the premise, to predict the support of a hypothesis. Verifying the entailment of such complex premises has important applications in rationale-generating LLMs, such as chain-of-thoughts (CoT) Wei et al. (2022), that typically generate multi-sentence rationales with incomplete information and suffer from inconsistent reasoning Ye and Durrett (2022). While EV is similar to Natural Language Inference (NLI) Dagan et al. (2006); Manning and MacCartney (2009), some interesting challenges distinguish it from NLI. Many existing textual inference datasets, such as SNLI Bowman et al. (2015), MNLI Williams et al. (2018), etc., mostly contain short sentence premises that only partially encapsulate the challenges of multi-sentence premises requiring complex reasoning. Predicting the entailment of such complex premise-hypothesis pairs often requires multi-hop reasoning, inferring missing information, etc., which is lacking from standard NLI datasets Gururangan et al. (2018); McCoy et al. (2019). Thus, studying EV in the context of modern LLMs is essential and is currently missing.
To this end, we first compile an evaluation benchmark to study the EV problem by selecting multiple datasets across three categories: NLI, contextual QA, and rationales. As shown in Table LABEL:tab:dataset_comparisons, the datasets typically contain multi-sentence premises that require inferring different types of knowledge to predict the entailment. Thus, this benchmark is more complex than standard NLI datasets and can be used as a new evaluation benchmark.
@@
Next, we evaluate LLMs and humans on this benchmark and make some interesting observations. Cognitive studies Buschman et al. (2011); Cowan (2001) have shown that an average human brain has a limited capacity to retain only four chunks in short-term memory, indicating a limitation of human inference abilities over long contexts. Our analysis shows that LLMs are indeed stronger than humans at tasks that involve multi-hop reasoning across long contexts, reinforcing this human limitation. In contrast, humans are better in cases that require simple deductive reasoning using substitutions, negations, etc., indicating that current LLMs lack consistency along these reasoning aspects. Further, humans and LLMs perform comparably in instances requiring inferring missing knowledge. These findings are depicted in Figure LABEL:fig:motivation with motivating examples.
@@
Additionally, on comparison between LLMs, we find that models finetuned on a specific dataset category are usually strong within the category but don’t generalize well on unseen categories. In contrast, instruction-finetuned models are better on average, with the best performing model, GPT-4, achieving 0.79 macro-F1 and outperforming the best-open-sourced model Flan-T5-xxl by 0.08 macro-F1 on average. In order to bridge this gap, we finetune a Flan-T5 Chung et al. (2022) model on a training subset containing datasets from each category of the above data collection. To this end, we explore two different training approaches: a classification-based finetuning that learns to directly predict the label and a ranking-based finetuning that learns to rank the most supported hypothesis from a given pair of hypotheses for a given premise. Ranking-based finetuning is better than classification, specifically in contextual QA datasets, as it can learn a softer decision boundary. Overall, our fine-tuned Flan-T5 outperforms GPT-3.5, baseline Flan-T5, and performs comparably to GPT-4 on the benchmark, thus providing a strong open-sourced model for entailment verification.
@@
Finally, we demonstrate the utility of our finetuned models on a downstream application of filtering unfaithful model-generated explanations. Self-consistency (SC) Wang et al. (2023) decoding samples multiple model-generated reasoning paths from the LLM decoder and aggregates them to predict the most consistent answer. We use our finetuned Flan-T5 to filter out non-entailed reasoning chains before aggregating the final prediction, which leads to 6% performance improvement on average across three MCQ datasets.
Overall, our contributions can be summarized as follows:
  1. 1.
    Using existing resources, we develop a benchmark for Entailment Verification (EV) of multi-sentence premises for both training and evaluation purposes.
  2. 2.
    We compare and contrast humans and GPT-4 on this benchmark and conclude that humans are more robust on simple deductive reasoning while LLMs excel at complex inferences.
  3. 3.
    We use a ranking-based objective to finetune Flan-T5-xxl model for the EV task, achieving comparable performance to GPT-4 and demonstrate the utility of our model in filtering non-entailed CoT rationales.
2 Entailment Verification using LLMs
@@
In this section, we formally define the task of entailment verification (EV), the datasets used to create the evaluation benchmark, and the evaluation procedure for evaluating different LLM baselines.
For a given premise $p$ and a hypothesis (or claim) $h$, the task of entailment verification is to determine whether the context has information that directly confirms the hypothesis or not, i.e., whether the hypothesis follows from the information present in the context. This is a binary classification task defined as $f(p,h)=\{\textrm{{support}},\textrm{{not support}}\}$, where $f$ is a classifier (human/LLM).
@@ 2.1 Evaluation Benchmark
In this section, we list some desirable properties we want to include in the EV benchmark, followed by details of the dataset categories we select.
@@
  • Type of Knowledge: Often, one or more information in the premise needs to be used to predict support. We categorize this information as entity-grounded, commonsense, or localized. Entity-grounded knowledge consists of information about entities and other general knowledge verifiable on the internet. These can be facts about general science, history, etc., or details of some known person, event, etc. It is possible to infer this information even if not mentioned in the premise. The commonsense knowledge is typically all information about everyday life that humans use implicitly but cannot always be verified online. This information is often missing from the premise and has to be inferred implicitly. Lastly, localized information is all other knowledge provided for understanding the events, people, or items mentioned in the premise that are not grounded to any known entity. This information depends on the premise’s specific context and, thus, is impossible to infer unless stated explicitly. Please refer to Table LABEL:tab:knowledge_examples for examples of each knowledge type.
@@
Natural Language Inference
Given the close connection between NLI and EV, it is an obvious choice to consider appropriate NLI datasets for the benchmark. To convert an NLI dataset for EV, we merge the neutral and contradict labels to the not support label. We use the following NLI datasets in our benchmark: WaNLI Liu et al. (2022), FEVER Nie et al. (2019), and ANLI Nie et al. (2020).
@@
Next, we consider multiple-choice question-answering datasets where the task is to answer a question based on a given context and some options. We use a QA-to-statement converter model Chen et al. (2021) to generate a hypothesis statement for each question option pair. Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support”. This process is depicted in Figure LABEL:fig:contextual_qa in Appendix A. We include the following datasets from this category: Cosmos QA (CosQA) Huang et al. (2019), SocialIQA (SIQA) Sap et al. (2019), DREAM Sun et al. (2019), BoolQ Clark et al. (2019), and RACE Lai et al. (2017).
@@
Lastly, we consider data sources where human-annotated explanations are available that justify the original hypothesis. In this case, we use the rationales as the premise. We use the following datasets: Entailer Tafjord et al. (2022), and ECQA Aggarwal et al. (2021).
2.2 Evaluation Metric
@@ 2.3 LLM Evaluation Setup
We evaluate two types of LLMs on the task of EV, as categorized below:
@@
In this, the models considered are already finetuned on some subset of the benchmark. We evaluate two models: RoBERTa Liu et al. (2019) (finetuned on NLI datasets) and Entailer-11B Tafjord et al. (2022) (finetuned on Entailer dataset). Refer to Appendix B.1 for more details on the evaluation setup for these models.
@@
Instruction-finetuned LLMs
These are language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions. Here, we evaluate Flan-T5-xxl Chung et al. (2022), GPT-3.5 Brown et al. (2020), and GPT-4 OpenAI (2023) models. To compute the label, we first modify a given premise-hypothesis pair $(p,h)$ into a prompted input $\mathcal{P}$ using the prompt template as shown in Box 1. Next, we compute a score $s$ as defined below:
where $p_{LLM}(\cdot|\mathcal{P})$ is the model’s probability distribution over the vocabulary. If the score $s$ is higher than a threshold (typically set to $0.5$ in all our experiments), we assign the label support, else we assign the label not support. For GPT-4 evaluation, we directly check for the “Yes” / “No” label prediction as the token probabilities are not accessible via the API. Please refer to Appendix B.2 for more details about the models and ablations on prompts.
3 Evaluation of Humans and LLMs
First, we randomly sample 100 instances from each dataset (i.e., 1000 instances in total) and conduct a human evaluation on this subset to estimate average human performance. Please refer to Appendix E.1 for more details on the annotation procedure. Additionally, we evaluate the above LLMs on this sampled subset and report those numbers for fair comparisons with humans. Table LABEL:tab:human_results shows the overall evaluation results.
3.1 Comparison among LLMs
We observe that the task-finetuned models (rows 1-2 in Table LABEL:tab:human_results) are weaker on average compared to the instruction-finetuned models (rows 3-5 in Table LABEL:tab:human_results). However, there are some interesting exceptions. RoBERTa, which is finetuned on NLI datasets, performs at par with GPT-4 on FEVER and outperforms it on WaNLI but falls behind on contextual QA and rationale datasets. On the other hand, Entailer-11B, which is finetuned on the Entailer dataset, outperforms GPT-4 on Entailer but lags on most of the other datasets. This demonstrates that finetuning an LLM using datasets from one of these categories is ineffective in outperforming models trained on more general data. Overall, we observe that GPT-4 is the best-performing entailment verification model on average and Flan-T5-xxl is the best open-source LLM for this task.
3.2 Comparison between Humans and LLMs
In Table LABEL:tab:human_results, we report the human performance and the corresponding annotation agreement [^3] We use a pairwise agreement ratio that computes the fraction of matched annotations over all pairs of annotations. (in parenthesis) for each dataset. We note that the agreement ratios are lowest for Contextual QA datasets, especially CosQA and SIQA. Questions in these datasets are often based on commonsense scenarios, and sometimes, the support for the wrong hypothesis can be debatable. Please refer to Appendix D for examples. Between humans and LLMs, we find that humans beat all the baseline LLMs, except GPT-4. This shows that existing open-sourced LLMs are subpar with humans on this task. Additionally, although humans and GPT-4 perform comparably on average, we note that large misalignments exist in different individual datasets. Specifically, we observe that ANLI, CosQA, SIQA, and Entailer are the four datasets with $>0.1$ absolute macro-F1 difference. We analyze these four datasets in Section 3.3.
@@ 3.2 Comparison between Humans and LLMs
3.3 Effect of Reasoning Type
We design an analysis to understand further the misalignments between GPT-4 and humans. First, we categorize the type of reasoning required to predict an entailment into the following four categories:
  • Simple Deductive (R1): The premise contains sentences that can be minimally combined in one step to predict the support for the hypothesis. This typically tests skills such as substitution, understanding negations, word meanings, etc.
  • Complex Deductive (R2): More than one step of reasoning is required to solve the task. Typically, this tests skills like mathematical reasoning, combining multiple information in context, etc.
  • Missing Entity-grounded/Commonsense Knowledge (R3): In this, some essential commonsense or entity-grounded knowledge is missing in the premise. Both humans and LLMs can implicitly invoke such information from the memory or the parametric knowledge obtained from pertaining, respectively.
  • Missing Localized Knowledge (R4): In this, information very specific to the premise is missing. Typically, this is information about the subjects in the context and is not grounded in any known entities. It is practically impossible for humans or LLMs to infer such information.
We note that these categories are mutually exclusive. [^4] Deductive reasoning implies that the premise has all the necessary information. Thus, any missing knowledge instance falls under inductive reasoning. Figure LABEL:fig:reasoning_analysis2 depicts the aggregated results. The top plot shows the percentage of each reasoning type among 400 samples from ANLI, CosQA, SIQA, and Entailer, and the bottom plot compares the human and GPT-4 macro-F1 scores. Please refer to Appendices E.2, E.3, and E.4 for details on the annotation setup, examples of each reasoning type, and detailed analysis, respectively.
The first type in Figure LABEL:fig:reasoning_analysis2 is simple deductive reasoning ($\sim$ 25% data). Here, humans perform better than GPT-4 by a small margin. Instances that require simple deductive reasoning usually use substitutions, negations, paraphrasing, etc., to prove entailment. We find that humans are more robust than GPT-4 in performing such simple deductive reasoning tasks, which is also observed in prior works Sanyal et al. (2022); Nguyen et al. (2023).
Next, we find that GPT-4 significantly outperforms the humans on complex reasoning that constitutes $\sim$ 20% data. This usually requires two skills: understanding multiple relevant information in the premise and combining them for reasoning. GPT-4 is likely a stronger context processor, especially for long premises, since it has been trained on long-context data sources OpenAI (2023). In contrast, cognitive studies Buschman et al. (2011); Cowan (2001) have shown that an average human brain can retain only four chunks in short-term memory, thus limiting the long-context processing abilities of humans.
Lastly, we observe that approximately 30% of the data has some missing entity-grounded or commonsense information while $\sim$ 25% of the data has missing localized information. To correctly predict entailment in such cases, a system needs to infer some missing grounded knowledge while not hallucinating specific localized information not mentioned in the premise. We find that both humans and models are comparable across the reasoning types R3 and R4, with R4 being more challenging. This shows that both models and humans tend to hallucinate missing localized information.